11 D X where X ~ unif{1, . . . ,0), θ e N. max(X1,-..,An) 2. Suppose that X1,... ,Xn Let T = X(n) =
Answer #1

here,

since, the sufficient and complete statistics X(n)_{} has the lebesgue pdf ,

n n\theta ^{-n}x^{n-1} I_{(0,\theta )}(x)

E(X_{n})=n\theta ^{-n}\int_{0}^{\infty }x^{n}dx=(n/n+1)\theta

hence , an unbiased estimator of \theta is (n+1)X_{(n)}/n

Which is UMVUE.

Suppose, that \gamma =g(\theta ) ,where g is differentiable function on (0,\infty )

An unbiased esrimator h(X_{(n)}) of \gamma must satisfy ,

\theta ^{n}g(\theta )=n\int_{0}^{\infty }h(x)x^{n-1}dx for all   \theta > 0

differentiating both size of previous equations and apply result of different of an integral lead to

n\theta ^{n-1}g(\theta )+\theta ^{n}g'(\theta )=nh(\theta )\theta ^{n-1}

Hence , the UMVUE of \gamma is

h(X_{n})=g(X_{(n)})+n^{-1}X_{(n)}g'(X_{(n)})

where, \gamma = real valued parameter related to population .

In particular , if \gamma =\theta ,then the UMVUE of \theta is (1+n^{-1})X_{(n)}